# Propagation of light in cold emitter ensembles with quantum position correlations due to static long-range dipolar interactions

### 20 Dec 2023 v2

@@ I Introduction
Disordered media of resonant light emitters provide rich mesoscopic systems [24, 25, 26] where light can mediate strong interactions. Cold and dense atomic ensembles in the limit of low light intensity (LLI) represent systems of dipolar emitters where light can induce strong position-dependent correlations, even though each individual atom’s response to a coherent light field behaves like a linear classical oscillator [27, 28, 29]. The light-mediated resonant DD interactions are highly sensitive to atomic positions, leading to radiative excitations in each atom that are correlated to the positions of every other atom in the sample. This granularity of atoms, which determines their optical response, cannot be captured by standard electrodynamics of continuous media, where light-induced interactions between atoms are treated in an averaged sense [29, 30, 31, 32]. In the limit where the atoms are considered stationary during imaging, the collective optical response of $N$ atoms is dependent upon the position correlation functions $\rho_{j}({\bf r}_{1},\ldots,{\bf r}_{j})$, for $j=1,\ldots,N$, that are determined before the light interacts with the sample [28]. When atomic positions are independent and uncorrelated, all these initial correlations factorize $\rho_{j}({\bf r}_{1},\ldots,{\bf r}_{j})=\rho_{1}({\bf r}_{1})\ldots\rho_{1}({\bf r}_{j})$, and solving for the optical response involves calculating the radiative excitations of the atoms that can, nevertheless, still be correlated due to the fluctuations of the atomic positions [33, 34]. However, highly non-trivial position correlation functions between electronic ground-state atoms $\rho_{j}({\bf r}_{1},\ldots,{\bf r}_{j})$ can exist due to quantum statistics and atomic interactions. Aside from the fundamental interest in cooperative responses in such systems, the scattered light then also directly conveys information about these correlations, serving as a diagnostic method.
@@
The strength of recurrent scattering in cooperative optical responses depends on the interparticle separation in units of the resonance wavevector of light. Increased density leads to more pronounced deviations from the single-atom Lorentzian lineshape. To characterize the impact of the static DD interactions, we systematically vary the strength of these interactions while approximately maintaining a constant peak density. Our observations reveal that, in both prolate and oblate traps, repulsive static interactions lead to short-range ordering among the dipoles, which, in turn, curtails fluctuations in the light-induced resonant DD interactions, particularly as these interactions become very large at short interparticle separations. This phenomenon is identified in increased coherent reflection and optical depth that are accompanied by reduced incoherent scattering. In an oblate trap, coherent transmission and reflection resonances narrow at low density but broaden at high density. In a prolate trap, the scattered light resonance can narrow even below the natural linewidth. The collective resonance shifts are substantially smaller than those predicted by the collective Lamb shift in continuous medium electrodynamics, underscoring the violation of the standard optics in the system [29, 30, 40, 41]. Intriguingly, especially in a prolate trap, the presence of the static DD interactions enables better targeted excitation of subradiant eigenmodes at high densities, although the linewidth of the eigenmodes can exhibit significant variation between different realizations due to the position fluctuations in the ensemble. Additionally, we find that the main difference between the two-level and the isotropic $J=0\rightarrow J^{\prime}=1$ transitions is the emergence of resonances where the orientations of the dipoles in an oblate trap, in the latter case, are parallel to the normal of the trap plane.
@@
The article is organized as follows: Section II provides the theoretical background for the static interactions and light-matter coupling, and Sec. III for stochastic electrodynamics and quantum Monte Carlo methods. The optical responses in an oblate trap, when the strength of the static interactions is varied for a constant peak density, are in Sec. IV.1, the differences between two-level and isotropic transitions in Sec. IV.2, inhomogeneous broadening in Sec. IV.3, and the results in a prolate trap in Sec. IV.4. The localization of excitations is presented in Sec. IV.5 and the optical responses for variable densities in Sec. IV.6 before concluding remarks in Sec. V.
@@ II.1 Static long-range dipole-dipole interactions
where $\bm{\mathbf{r}}_{\ell j}=\bm{\mathbf{r}}_{j}-\bm{\mathbf{r}}_{\ell}$ is the vector joining the atoms $j$ and $\ell$. For magnetic dipoles, we have $C=\mu_{0}$. The DD interaction can be characterized by the interaction length
@@ II.2 Collective atom-light interactions
We solve the optical response of the atoms in the limit of LLI where the individual atoms respond to light linearly and behave as a set of coupled, driven harmonic oscillators. In this section, we first consider commonly employed basic formalism [42] for the coupled dynamics between the atoms and light for a fixed set of atomic positions $\{\textbf{r}_{1},\textbf{r}_{2},\ldots,\textbf{r}_{N}\}$. When the positions fluctuate, these equations are then utilized in stochastic electrodynamics simulations for a disordered system in Sec. III.1.
@@ III.1 Stochastic electrodynamics
The ground-state correlations $\rho_{k}$ exist before the light enters the sample. ${\bf P}_{p}({\bf r}_{1},\ldots,{\bf r}_{p-1};{\bf r}_{p})$ represents the correlation function for ground state atoms at ${\bf r}_{1},\ldots,{\bf r}_{p-1}$, given that there is polarization at ${\bf r}_{p}$. The strong coupling emerges from the third term on the right-hand side, representing repeated exchanges of a photon between the same atoms, giving rise to recurrent scattering processes.
@@ III.3 Scattered light
Since we consider here the LLI limit where each atom responds to light classically, the scattered light from the atomic ensemble is directly evaluated from the stochastic simulation data by ensemble-averaging over many realizations. We then have for the transmission $T$ and reflection $R$
@@ IV.1 Varying static dipolar interaction in an oblate trap
The crucial parameter in the cooperative response of independent atoms is the atom separation in the units of $1/k$ that determines the strength of recurrent scattering and the emergence of light-induced correlations (Sec. III.1[27, 30]. Increasing densities lead to deviations from the continuous media electrodynamics due to these correlations. To isolate the effect of static DD interactions and the resulting position distributions on the optical response, we therefore consider atomic ensembles with constant ${{\bar{\rho}}_{\rm{2D}}/k^{2}}$ while varying ${R_{\rm{dip}}}$ [Eq. (4)]. We express the interaction strength as the ratio between the dipolar length ${R_{\rm{dip}}}$ [Eq. (4)] and the average separation between atoms in the $xy$ plane at the trap center ${R_{\rm{dip}}\bar{\rho}_{2\rm{D}}^{1/2}}$ (in a prolate trap we use ${R_{\rm{dip}}\bar{\rho}_{1\rm{D}}}$). We take $N=200$ atoms and adjust the peak 2D density ${{\bar{\rho}}_{\rm{2D}}}$ by changing the trapping frequencies. The transmission and reflection are shown in Fig. 4 for two-level atoms for linear polarization in the trap plane at the densities ${{\bar{\rho}}_{\rm{2D}}/k^{2}\simeq 0.1}$ and 1, respectively, by varying ${R_{\rm{dip}}\bar{\rho}_{2\rm{D}}^{1/2}\simeq 0,0.017,0.15,1.6}$.
@@
The effect of the static DD interactions on the light-mediated interactions between the atoms at a given density is due to position correlations. Increasing the repulsive interactions modify the distribution of interatomic separations introducing short-range ordering of the atoms, as shown in the probability distributions of the nearest-neighbor and all-atom separations in Figs. 6 and 7. As ${R_{\rm{dip}}}$ and repulsion increase, the atoms are prevented from coming close to each other. This suppression of small interatomic separations prevents the light-mediated DD interactions, which scale as $\propto 1/r^{3}$ at small separations, becoming very large. In disordered ensembles this reduces the fluctuations of light-induced resonance shifts between any atom pairs, resulting in reduced inhomogeneous broadening of resonance frequencies and incoherent scattering. In Fig. 7, the most pronounced short-range ordering occurs for ${R_{\rm{dip}}\sqrt{{\bar{\rho}}_{\rm{2D}}}\simeq 1.6}$ when the atoms are unable to approach each other, representing the most dramatic lineshape deformation in Fig. 4. We additionally find in Figs. 6 and 7 that the large separations are suppressed due to increased trap frequencies that maintain the constant peak density value.
@@ IV.4 Varying static dipolar interaction in a prolate trap
Figure 13: Scattered light resonance (a) widths and (b) shifts of maxima from Fig. 12 for ${\bar{\rho}_{\rm{1D}}/k\simeq 1}$. HWHM is calculated by averaging the contributions of the left and right sides of the peak.

### 11 Jan 2024 v3

@@ I Introduction
Disordered media of resonant light emitters provide rich mesoscopic systems [24, 25, 26] where light can mediate strong interactions. Cold and dense atomic ensembles in the limit of low light intensity (LLI) represent systems of dipolar emitters where light can induce strong position-dependent correlations, even though each individual atom’s response to a coherent light field behaves like a linear classical oscillator [27, 28, 29]. The light-mediated resonant DD interactions are highly sensitive to atomic positions, leading to radiative excitations in each atom that are correlated to the positions of every other atom in the sample. This granularity of atoms, which determines their optical response, cannot be captured by standard electrodynamics of continuous media, where light-induced interactions between atoms are treated in an averaged sense [29, 30, 31, 32]. In the limit where the atoms are considered stationary during imaging, the collective optical response of $N$ atoms is dependent upon the position correlation functions $\rho_{j}({\bf r}_{1},\ldots,{\bf r}_{j})$, for $j=1,\ldots,N$, that are determined before the light interacts with the sample [28]. When atomic positions are independent and uncorrelated, all these initial correlations factorize $\rho_{j}({\bf r}_{1},\ldots,{\bf r}_{j})=\rho_{1}({\bf r}_{1})\ldots\rho_{1}({\bf r}_{j})$, and solving for the optical response involves calculating the radiative excitations of the atoms that can, nevertheless, still be correlated in a disordered system [33, 34]. However, highly non-trivial position correlation functions between electronic ground-state atoms $\rho_{j}({\bf r}_{1},\ldots,{\bf r}_{j})$ can exist due to quantum statistics and atomic interactions. Aside from the fundamental interest in cooperative responses in such systems, the scattered light then also directly conveys information about these correlations, serving as a diagnostic method.
@@
The strength of recurrent scattering in cooperative optical responses depends on the interparticle separation in units of the resonance wavenumber of light. Increased density leads to more pronounced deviations from the single-atom Lorentzian lineshape. To characterize the impact of the static DD interactions, we systematically vary the strength of these interactions while approximately maintaining a constant peak density. Our observations reveal that, in both prolate and oblate traps, repulsive static interactions lead to short-range ordering among the dipoles, which, in turn, curtails fluctuations in the light-induced resonant DD interactions, particularly as these interactions become very large at short interparticle separations. This phenomenon is identified in increased coherent reflection and optical depth that are accompanied by reduced incoherent scattering. In an oblate trap, coherent transmission and reflection resonances narrow at low density but broaden at high density. In a prolate trap, the scattered light resonance can narrow even below the natural linewidth. The collective resonance shifts are substantially smaller than those predicted by the collective Lamb shift in continuous medium electrodynamics, underscoring the violation of the standard optics in the system [29, 30, 40, 41]. Intriguingly, especially in a prolate trap, the presence of the static DD interactions enables better targeted excitation of subradiant eigenmodes at high densities, although the linewidth of the eigenmodes can exhibit significant variation between different realizations due to the position fluctuations in the ensemble. Additionally, we find that the main difference between the two-level and the isotropic $J=0\rightarrow J^{\prime}=1$ transitions is the emergence of resonances where the orientations of the dipoles in an oblate trap, in the latter case, are parallel to the normal of the trap plane.
@@
The article is organized as follows: Section II provides the theoretical background for describing static interactions and light-matter coupling, while Sec. III gives an overview of stochastic electrodynamics and the quantum Monte Carlo methods. The optical responses in an oblate trap, when the strength of the static interactions is varied for a constant peak density, are in Sec. IV.1, the differences between two-level and isotropic transitions in Sec. IV.2, inhomogeneous broadening in Sec. IV.3, and the results in a prolate trap in Sec. IV.4. The localization of excitations is presented in Sec. IV.5 and the optical responses for variable densities in Sec. IV.6 before concluding remarks in Sec. V.
@@ II.1 Static long-range dipole-dipole interactions
where $\bm{\mathbf{r}}_{\ell j}=\bm{\mathbf{r}}_{j}-\bm{\mathbf{r}}_{\ell}$ is the vector joining the atoms $j$ and $\ell$ and $\hat{{\bf r}}_{\ell j}={\bf r}_{\ell j}/|{\bf r}_{\ell j}|$. For magnetic dipoles, we have $C=\mu_{0}$. The DD interaction can be characterized by the interaction length
@@ II.2 Collective atom-light interactions
We solve the optical response of the atoms in the limit of LLI where the individual atoms respond to light linearly and behave as a set of coupled, driven harmonic oscillators. In this section, we first consider the commonly employed basic formalism [42] for the coupled dynamics between the atoms and light for a fixed set of atomic positions $\{\textbf{r}_{1},\textbf{r}_{2},\ldots,\textbf{r}_{N}\}$. When the positions fluctuate, these equations are then utilized in stochastic electrodynamics simulations for a disordered system in Sec. III.1.
@@ III.1 Stochastic electrodynamics
The ground-state correlations $\rho_{p}$ exist before the light enters the sample. ${\bf P}_{p}({\bf r}_{1},\ldots,{\bf r}_{p-1};{\bf r}_{p})$ represents the correlation function for ground state atoms at ${\bf r}_{1},\ldots,{\bf r}_{p-1}$, given that there is polarization at ${\bf r}_{p}$. The strong coupling emerges from the third term on the right-hand side, representing repeated exchanges of a photon between the same atoms, giving rise to recurrent scattering processes.
@@ III.3 Scattered light
Since we consider here the LLI limit where each atom responds to light classically, the scattered light from the atomic ensemble is directly evaluated from the stochastic simulation data by ensemble-averaging over many realizations [55]. We then have for the transmission $T$ and reflection $R$
@@ IV.1 Varying static dipolar interaction in an oblate trap
The crucial parameter in the cooperative response of independent atoms is the atom separation in the units of $1/k$ that determines the strength of recurrent scattering and the emergence of light-induced correlations (Sec. III.1[27, 30]. Increasing densities lead to deviations from the continuous media electrodynamics due to these correlations. To isolate the effect of static DD interactions and the resulting position distributions on the optical response, we therefore consider atomic ensembles with constant ${{\bar{\rho}}_{\rm{2D}}/k^{2}}$ while varying the dipolar length ${R_{\rm{dip}}}$ [Eq. (4)]. We express the interaction strength as the ratio between ${R_{\rm{dip}}}$ and the average separation between atoms in the $xy$ plane at the trap center ${R_{\rm{dip}}\bar{\rho}_{2\rm{D}}^{1/2}}$ (in a prolate trap we use ${R_{\rm{dip}}\bar{\rho}_{1\rm{D}}}$). We take $N=200$ atoms and adjust the peak 2D density ${{\bar{\rho}}_{\rm{2D}}}$ by changing the trapping frequencies. The transmission and reflection are shown in Fig. 4 for two-level atoms for linear polarization in the trap plane at the densities ${{\bar{\rho}}_{\rm{2D}}/k^{2}\simeq 0.1}$ and 1, respectively, by varying ${R_{\rm{dip}}\bar{\rho}_{2\rm{D}}^{1/2}\simeq 0,0.017,0.15,1.6}$.
@@
The effect of the static DD interactions on the light-mediated interactions between the atoms at a given density is due to position correlations. Increasing the repulsive interactions modify the distribution of interatomic separations introducing short-range ordering of the atoms, as shown in the probability distributions of the nearest-neighbor and all-atom separations in Figs. 6 and 7. As ${R_{\rm{dip}}}$ and repulsion increase, the atoms are prevented from coming close to each other. This suppression of small interatomic separations prevents the light-mediated DD interactions, which scale as $\propto 1/r^{3}$ at small separations, becoming very large. In disordered ensembles this reduces the fluctuations of light-induced resonance shifts between any atom pairs, resulting in reduced inhomogeneous broadening of resonance frequencies due to the resonance shifts and reduced incoherent scattering. In Fig. 7, the most pronounced short-range ordering occurs for ${R_{\rm{dip}}\sqrt{{\bar{\rho}}_{\rm{2D}}}\simeq 1.6}$ when the atoms are unable to approach each other, representing the most dramatic lineshape deformation in Fig. 4. We additionally find in Figs. 6 and 7 that the large separations are suppressed due to increased trap frequencies that maintain the constant peak density value.
@@ IV.4 Varying static dipolar interaction in a prolate trap
Figure 13: Scattered light resonance (a) widths and (b) shifts of maxima from Fig. 12 for ${\bar{\rho}_{\rm{1D}}/k\simeq 1}$. The HWHM is calculated by averaging the contributions of the left and right sides of the peak.